GARR Integrated Networking Suite
User: Password: SSO login

Ticket info

Ticket number 26784
Context NOC-NetworkServices
Ticket Status Close
Ticket Source GARR
Ticket Type Maintenance
Ticket open time 12/01/2023 12:33
Ticket close time 27/01/2023 09:47
Affected Link e2e LHCOPN CERN backup
Affected Service e2e LHCOPN MILANO-CERN
Problem type Fault
Problem description Planned maintenance
Problem fixer Supplier
Problem start 16/02/2023 21:00
Problem end 17/02/2023 05:00
History 12/01/2023 12:33
The circuit provider has announced a planned maintenance in the above window to perform a replacement of a fibre cable. The entire window is reserved for outage events.
Problem fix maintenance has now been completed